# Re(l)ality:

### v3

@@ Abstract
Using the fiber bundle framework, this work investigates the conceptual and mathematical foundations of reference frames in General Relativity by contrasting two paradigms. The View from Nowhere interprets frame representations as perspectives on an invariant equivalence class, while the View from Everywhere posits each frame representation as constituting a fully-fledged reality itself. What emerges is a conception of reality that I term ”Relality.” The paper critically examines the philosophical and practical implications of these views, with a focus on reconciling theory with experimental practice. Central to the discussion is the challenge of providing a perspicuous characterisation of ontology. The View from Nowhere aligns with the so-called ‘ sophisticated approach to symmetries’ and complicates the empirical grounding of theoretical constructs. In contrast, the View from Everywhere offers a relational ontology that avoids the abstraction of equivalence classes.
@@ 1 Introduction
All measurements are inherently local, as they rely on observations within specific regions of spacetime. Much of our experience involves measuring locally defined variables at particular spacetime points or within localized regions. As Gary and Giddings (2007) emphasise, “all we can truly observe is localised — we have no access to infinity.” In GR, gauge symmetries are expressed as diffeomorphisms. As a result, any quantity defined locally in spacetime does not remain invariant under these transformations, raising challenges for constructing local, gauge-invariant observables.
@@
This relational framework is closely connected to the role of reference frames in GR, as discussed by Bamonti (2023). Reference frames offer a structured way to define gauge-invariant quantities relationally. In practice, constructing relational observables often involves treating one partial observable as a spatiotemporal reference frame for another, thereby highlighting the inherently relational nature of locality in this framework.
@@
The first, the, View from Nowhere is the dominant perspective in the literature. It asserts that changing a reference frame does not affect the physical content of the description, as a reference frame provides merely a perspective on a shared, objective reality. This idea has its roots in Special Relativity, where different inertial reference frames offer distinct perspectives on a Lorentz-invariant physical reality. Prominent examples of this view include the perspective-neutral framework introduced by Vanrietvelde et al. (2020) and related work by Giacomini et al. (2019) and Kabel et al. (2024). Importantly, I will use the term ‘perspective’ in the ordinary deflated sense throughout the paper. [^4] The term ’perspective’ is widely used in the philosophy of science. For example, Michela Massimi uses the notion of perspective in terms of a historically and culturally situated epistemic lens within a debate on scientific realism Massimi (2022).
@@
The View from Everywhere, proposed in this paper, challenges the assumption of an underlying objective reality. I argue that there is no direct analogy between GR and Special Relativity in this context. While Special Relativity relies on a fixed, Lorentz-invariant framework, GR introduces local gauge-invariant quantities that are not necessarily frame-invariant. Local relational observables, although gauge-invariant, can differ numerically when expressed in different reference frames. Consequently, the View from Everywhere suggests that frame representations are not merely perspectives on an objective reality; they are fundamental realities themselves.
@@
In the context of non-relational quantities, we can distinguish quantities that are:
@@
    Spatiotemporally explicit: expressible by a single model as explicit functions of points of $\mathcal{M}$. They can be distinguished into:
@@
      Spatiotemporally local: locally expressible by a single model, e.g. the metric tensor on the manifold $g_{ab}(p)$. Such objects are not gauge-invariant
@@
      Spatiotemporally non-local: non-locally expressible by a single model, e.g. the total volume integral of a given metric. Such objects are gauge-invariant
@@
    Spatiotemporally implicit: not expressible by a unique spatiotemporal model, but only as equivalence classes. They are usually denoted using square brackets: for example, $[g_{ab}]$. Such objects are gauge-invariant.
Once reference frames are introduced, it is straightforward to construct relational quantities. Analogously as done above, we can distinguish various possible formulations of relational objects in GR, when written in terms of reference frames. Here, I provisionally denote a reference frame without further specification as a set of 4 scalar quantities $\phi^{(I)}\mid_{I=1,\dots,4}$ associating a real quadruple with each point $P$ of the manifold. (See the next section for a more detailed discussion). We can distinguish between relational quantities that are:
@@
    Frame explicit: expressible by a single model as explicit functions of a reference frame. They can be distinguished into:
@@
      Frame local: locally expressible by a single model $g_{IJ}(\phi)$ (see the next section for the detailed construction and meaning of symbols). Such objects are gauge-invariant. However, using uncoupled reference frames they could also be only relational, but not gauge-invariant (see Bamonti and Gomes (2024b)).
@@
      Frame non-local: non-locally expressible by a single model, e.g. the total volume integral of a given gauge-invariant metric $g_{IJ}$. Such objects are always gauge-invariant, independently if the reference frame is coupled with the metric or not
@@ 2. II)
    Frame implicit: not expressible by a unique spatiotemporal model, but only as equivalence classes. In such a case, the equivalence class is the collection of all spatiotemporally explicit frame-explicit models. They can be denoted using square brackets: for example $[g_{IJ}]$, which constitutes a gauge-invariant object. See section 4 for a discussion on such object.
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality typical of the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^6] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happening of events, which specify the ”when” and ”where” of physical phenomena in a relational manner. Accordingly, the ‘where and when’ are consequential to the happening of an event (which is a gauge-invariant observable). So it is the happening of the event that determines where and when it happens and not the event happening at a where and when. In this paper I choose the more conservative option and identify space-time with option (ii).
@@
    Reference frame-dependence: a quantity is reference frame-dependent if its definition depend on the reference frame in which is defined
@@
    Reference frame-independence: a quantity is reference frame-independent if its definition does not depend on the reference frame in which it is, however, defined
@@
2 The Bundle Formalism: A Gauge Perspective on Reference Frames
In this section, I use the fiber-bundle formalism to describe reference frames and relational observables. This approach, widely used in foundational studies of gauge theories (see e.g. Healey (2007); Weatherall (2016)), is primarily attributed to the work of Gomes (see, e.g., Gomes (2023a, b) for a rigorous treatment). Let $M$ represent the space of models $m$ of the theory. The space $M$ can be described as a principal bundle with $S$ as its structure group and $[M]:=\{[m]\mid m\in M\}$ as its base space, where $[m]$ identifies the equivalence class of models under the transformations of $\mathcal{S}$. In this formalism, selecting a reference frame involves defining a unique section map $\sigma:[m]\to\sigma([m])\in M$, which smoothly maps equivalence classes of models to individual models in the space . This corresponds to choosing a submanifold in the fiber bundle that intersects each fiber $\mathbb{F}_{m}:=\mathrm{pr}^{-1}([m])$ exactly once, with $\mathrm{pr}:m\to[m]$ being the projection map.
Choosing a reference frame can also be interpreted as selecting a specific gauge. In fact, the choice of a gauge is equivalent to the choice of a reference frame (see Dittrich (2007) for insights on the relationship between gauge-fixed observables and relational observables). In gauge theories with a principal fiber bundle structure, gauge-fixing determines a section through the fiber bundle. Given a symmetry group $S$, each fiber corresponds to a gauge orbit — the set of all configurations of a field that are related by gauge transformations, generated by the constraints of the theory. Specifically, there exists a one-to-one correspondence between each gauge orbit and the equivalence class of models under the symmetry $s\in S$. More concretely, for $m\in M$ and $s\in S$, the orbit $\mathcal{O}_{m}$ is defined as $\mathcal{O}_{m}=\{m_{s}\mid s\in S\}$, where $m_{s}$ denotes the model to which the transformation $s$ is applied. Due to the symmetry of the theory, the model space includes significant redundancy. The physical content of the theory is captured by a single representative from each gauge orbit, while other models within the same orbit are redundant, isomorphic copies. [^7] This does not imply that gauge freedom constitutes mere ”descriptive fluff” (see Earman (2004)). On the contrary, the correspondence between gauge-fixing and reference frame selection highlights the relational nature of physics. The additional degrees of freedom are meaningful as they represent the possible ways a system can form observables relative to another system. Isomorphic models provide ”handles” through which systems can couple; see Rovelli (2014) and Adlam (2024). This redundancy is typically resolved by fixing a gauge. Similarly, in GR, one fixes a reference frame through a coordinate gauge condition (see Bamonti (2023); Gomes (2024b)).
@@ 2 The Bundle Formalism: A Gauge Perspective on Reference Frames
where $f_{\Re_{g}}$ acts as the equivalent of a section map within the fiber-bundle framework. Specifically, this diffeomorphism $f_{\Re_{g}}:g_{ab}\to f_{\Re_{g}}^{*}g_{ab}$ serves as a projection operator, uniquely mapping any element of a given fiber to the section’s image. It is also called the projection operator for the section. [^13] The projection corresponding to the choice of a reference frame can be written down in terms of coordinate charts. For instance, in GR two common gauge-fixings are the De Donder gauge in the Lagrangian sector which corresponds to the condition $F(g)=\partial_{\mu}(g^{\mu\nu}\sqrt{g})=0$, and the CMC gauge in the Hamiltonian sector: $F(h^{ij},\pi_{ij})=h^{ij}\pi_{ij}=const$. The former corresponds to the use of coordinates that satisfy a relativistic wave equation $\Box x^{\mu}=0$. The latter, selects global simultaneity homogeneous 3-hypersurfaces $\Sigma_{\tau}$, parametrised by a universal time $\tau$. Notice that both are only partial gauge-fixings. More precisely, $f_{\Re_{g}}$ is the embedding map, acting within a fibre, from the fibre bundle manifold of models $Lor(\mathcal{M})$ to the image of the section map, and it is characterised by the auxiliary condition $F_{\Re_{g}}(g_{ab})=0$. The constraint $F_{\Re_{g}}(g_{ab})=0$ defines a ‘level surface’ of the section map along the fibers, effectively making the choice of a reference frame (or section) analogous to a gauge-fixing procedure. See Figure 1.141414 According to the condition (3), the choice of a reference frame, seen as the choice of a section, is based on the imposition of a set of conditions that a models must satisfy. This procedure is analogous to what Gauss (1902) proposed in order to describe a surface embedded in an ambient space not from an external point of view, i.e. using the coordinates of the ambient space, but ‘standing on the surface itself’. Such an embedded surface is intrinsically describable using some parametric equations. For example, a generic $n$-dimensional surface $\Sigma\subset N$, embedded in a generic $(n+1)$-dimensional Euclidean ambient space $N$ (characterised by $n+1$ coordinates $x^{i}$), can be described by $n$ parameters $u^{\alpha}$ with parametric equations of the type: $x^{i}(u^{\alpha})=0$. In our case, such equations take the form $F_{\mathcal{R}_{g}}(g_{ab})=0$, with $F_{\mathcal{R}_{g}}\in C^{\infty}(\mathcal{M})$ being a smooth and regular function.  [^15] Given a general symmetry group, its gauge orbits are in general not one-dimensional. The representation of figure 1 is faithful only for one-dimensional groups, whose action can be depicted in a one-to-one manner along the one-dimensional orbits.

### v4

@@ Abstract
Using the fibre bundle framework, this work investigates the conceptual and mathematical foundations of reference frames in General Relativity by contrasting two paradigms. The View from Nowhere interprets frame representations as perspectives on an invariant equivalence class, while the View from Everywhere posits each frame representation as constituting a fully-fledged reality itself. What emerges is a conception of reality that I term ”Relality.” The paper critically examines the philosophical and practical implications of these views, with a focus on reconciling theory with experimental practice. Central to the discussion is the challenge of providing a perspicuous characterisation of ontology. The View from Nowhere aligns with the so-called ‘ sophisticated approach to symmetries’ and complicates the empirical grounding of theoretical constructs. In contrast, the View from Everywhere offers a relational ontology that avoids the abstraction of equivalence classes.
@@ 1 Introduction
All measurements are inherently local, as they rely on observations within specific regions of spacetime. Much of our experience involves measuring locally defined variables at particular spacetime points or within localised regions. As Gary and Giddings (2007) emphasise, “all we can truly observe is localised — we have no access to infinity.” In GR, gauge symmetries are expressed as diffeomorphisms. As a result, any quantity defined locally in spacetime does not remain invariant under these transformations, raising challenges for constructing local, gauge-invariant observables.
@@
This relational framework is closely connected to the role of reference frames in GR, as discussed in Bamonti (2023). Reference frames offer a structured way to define gauge-invariant quantities relationally. In practice, constructing relational observables often involves treating one partial observable as a spatiotemporal reference frame for another, thereby highlighting the inherently relational nature of locality in this framework.
@@
The first, the View from Nowhere is the dominant perspective in the literature. It often asserts that changing a reference frame does not affect the physical content of the description, as a reference frame provides merely a perspective on a shared, objective reality. This idea has its roots in Special Relativity, where different inertial reference frames offer distinct perspectives on a Lorentz-invariant physical reality. Prominent examples of this view include the perspective-neutral framework introduced by Vanrietvelde et al. (2020) and related work by Giacomini et al. (2019) and Kabel et al. (2024). Explicitly, Kabel et al. (2024) state “if we assume that the covariance of physical laws under classical changes of reference frame extends linearly to encompass changes of QRF, then the latter should not affect the physical situation. Thus, whether a system is in superposition or is entangled, and how these properties change dynamically, becomes a mere matter of perspective” [my italics].
Importantly, throughout the paper I will use the term ‘perspective’ in the sense also adopted in (Adlam, 2024b) and there called ‘physical perspectivalism’: a perspective is defined by physical observers which play the role of a reference frame. [^4] The term ’perspective’ is widely used in the philosophy of science. For example, Michela Massimi uses the notion of perspective in terms of a historically and culturally situated epistemic lens within a debate on scientific realism Massimi (2022). Another possibility is to use the term perspective only in an epistemic way: a perspective merely indicates the degree of partial knowledge of a conscious observer. This form of perspectivalism is known as ‘epistemic perspectivalism’.
It is worth clarifying the connection I am drawing between the View from Nowhere and Perspectivalism. In the literature, e.g. in Adlam (2024b), it is argued that perspectivalism rejects the View from Nowhere and that “perspectivalists are fond of the mantra ‘there is no view from nowhere’ ” (ivi, p.4). [^5] See also Ruyant (2020): “(various different forms of perspectivalism) share the general idea that there is no ‘view from nowhere’ ”. However, I believe this claim stems from a misuse of the term ‘perspective’. In particular, I argue that a perspective is always a perspective on something which stays unchanged. The perspective neutral structure presented in Vanrietvelde et al. (2020) is exactly this ‘something’ on which various perspectives are defined.
In characterising physical perspectivalism, Adlam advocates a moderate version, whereby perspective-neutral physical facts exist, but have no empirical meaning. Well, moderate perspectivalism advocates the existence of a View from Nowhere, only it does not give it empirical meaning, but it is there.666‘There’ where? I claim nowhere! See fn.22 in §3. As Adlam herself also claims:
It appears that considerations from quantum mechanics and relativity do provide support for the idea that empirically meaningful descriptions typically have to be relativized to a perspective. But this in and of itself does not tell us whether we should adopt mod- erate or strong physical perspectivalism. To make that choice, we must decide whether these considerations suggest that all facts about physical reality must be relativized to a perspective, or whether they instead support the existence of some perspective-neutral facts about physical reality, over and above descriptions relativized to perspectives (Adlam, 2024b, p.6).
In short, moderate perspectivalism understands each perspective as a partial, limited glimpse into a more complete reality, as is also evident from Adlam’s admission that moderate perspectivalism is motivated by the epistemic humility sustained by epistemic perspectivalists. A moderate perspectivalist can thus be seen as the proponent of a vision of a kind of ‘ontological humility’.
The strong version of perspectivalism, on the other hand, rejects the existence of any perspective-neutral fact: there are only perspectives. Based on what I argued earlier, this version of perspectivalism is a non-sense by definition: a perspective is always a perspective on something. Consequently, the perspectivalism that will be criticised in this paper in association with the View from Nowhere is the moderate version: the only one that makes sense.777 As will become clear later in the introduction, another critical point of Adlam’s cited work is that she does not distinguish between independence and freedom, but only adopt the quite ambiguous term ‘neutral’ and this leads her to consider that the presence of connections between perspectives and strong perspectivalism cannot be compatible. However, if the connections are considered perspective-independent facts, there is no incompatibility.
@@
The View from Everywhere, proposed in this paper, challenges the assumption of an underlying objective reality. I argue that there is no direct analogy between GR and Special Relativity in this context. While Special Relativity relies on a fixed, Lorentz-invariant framework which proves us with a “gods-eye’ description of the way in which all of the perspectives are related” (van Fraassen, 2008), GR introduces local gauge-invariant quantities that are not necessarily frame-invariant. Local relational observables, although gauge-invariant, can differ numerically when expressed in different reference frames. Consequently, the View from Everywhere suggests that frame representations are not merely perspectives on an objective reality; they are fundamental realities themselves. No God’s view is allowed.
@@
Within the framework of non-relational quantities, it is possible to differentiate between quantities that are:
@@
    Spatiotemporally explicit: A quantity is said to be spatiotemporally explicit if it is defined by a single function of a given region $U\subseteq\mathcal{M}$. They can be distinguished into:
@@
      Spatiotemporally local: A spatiotemporally local quantity is one that is determined solely by the properties at a single point, understood as a region of $\mathcal{M}$ having a zero Lebesgue measure. It is “pointwise” defined and does not depend on information from neighboring points. For example, the metric tensor $g_{ab}(p)$ is defined locally on the manifold—it provides the geometric structure exactly at $p$. Such objects are not gauge-invariant.
@@
      Spatiotemporally non-local: A spatiotemporally non-local quantity, on the other hand, is defined over an extended region of $\mathcal{M}$. Such quantities involve integrals or sums over areas or volumes, meaning that their values cannot be determined solely by the properties at an isolated point. For instance, the total volume obtained by integrating the metric over a region is non-local because it aggregates information from an extended domain of $\mathcal{M}$. Such objects are gauge-invariant. [^9] Strictly speaking, these quantities are gauge-invariant if the integration extremes are not fixed. Otherwise, also a volume varies even under a diffeomorphic transformation. See, e.g., formulas 15 and 17 in Rovelli (2014).
@@
    Spatiotemporally implicit: A quantity is spatiotemporally implicit if it cannot be represented by a single unique function on $\mathcal{M}$, but only as an equivalence class. Such quantity is usually denoted using square brackets: for example, $[g_{ab}]$. Such objects are gauge-invariant.
Introducing reference frames allows us to construct relational quantities. Here, I provisionally denote a reference frame without further specification as a set of four scalar quantities $\phi^{(I)}\mid_{I=1,\dots,4}$ uniquely associating a real quadruple $x\in\mathbb{R}^{4}$ with each point $p\in\mathcal{M}$ (see the next section for a more detailed discussion; see Bamonti (2023) for a comprehensive discussion). Similarly to the above, it is possible to distinguish different definitions of relational objects in GR, when written in terms of reference frames.
We can distinguish between relational quantities that are:
@@
    Frame explicit: These quantities are described by a single function of a given region ‘covered’ by the reference frame field (we can safely imagine a reference frame exactly as ‘chart’ in differential geometry). They can be subdivided into:
@@
      Frame local: A frame local quantity is one that is determined solely by the properties at a single point $p:=\phi^{-1}(x)$ – for example, $g_{IJ}(\phi)$ (see the next section for the detailed construction and meaning of notation). [^10] One might wonder whether $g_{IJ}(\phi)$ is not a functional, since each $\phi$ is a scalar field (a functions of point $p\in\mathcal{M}$). However, even though $\phi^{(I)}(p)$ are scalar functions defined on the manifold, when we write $g_{IJ}(\phi)$ we are expressing the metric components in terms of the values provided by each $\phi$ at each point. In other words, each point $p$ is mapped to a coordinate $x=\phi(p)\in\mathbb{R}^{4}$, and $g$ is a function that assigns the corresponding metric tensor $g_{IJ}$ to each such $x$. This is different from a functional, which would take an entire function as input and yield a number. Thus, $g_{IJ}(\phi)$ is a function, not a functional. Its value at a point $p$ depends on the local value of the reference fields at that point, so a frame-local quantity is defined by ‘pointwise’ information alone, where the notion of locality is relational in terms of other fields and not manifold points. These objects are gauge-invariant when the reference frame is appropriately coupled to the metric. However, if the reference frame is uncoupled, the quantity may remain relational without being fully gauge-invariant (see Bamonti and Gomes (2024b)).
@@
      Frame non-local: A frame non-local quantity is defined over an extended region parametrised by the reference frame. For example, total volume integral of a given gauge-invariant metric $g_{IJ}(\phi)$ locally defined is a frame non-local quantity, since its value depends on the metric values over a range of points rather than at a single point. Also such objects are gauge-invariant only if the reference frame is coupled with the metric. Otherwise, we can independently apply a diffeomorphism either to the integration extremes or to the metric and the situation is analogous to that in which after a diffeomorphism the metric (the integrand) changes, but the extremes stay fixed, resulting in the integral value changing (see fn.9 above).
@@ 2. II)
    Frame implicit: A frame implicit quantity is not expressible by a unique function, but only as a collection of frame-explicit quantities: for example an equivalence class. In such a case, the equivalence class can still be denoted using square brackets as $[g_{IJ}]$ and constitutes a gauge-invariant object. See §4 for a discussion on such object.
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality inherent to the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^11] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happening of events, which specify the ”when” and ”where” of physical phenomena in a relational manner. Accordingly, the ‘where and when’ are consequential to the happening of an event (which is a gauge-invariant observable). So it is the happening of the event that determines where and when it happens and not the event happening at a where and when. In this paper I choose the more conservative option and identify space-time with option (ii).
@@
    Reference frame-dependence: a quantity is reference frame-dependent if its definition depend on the reference frame in which is defined.
@@
    Reference frame-independence: a quantity is reference frame-independent if its definition does not depend on the reference frame in which it is, however, defined.
@@
2 The Fibre Bundle Formalism: A Gauge Perspective on Reference Frames
In this section, I use the fibre-bundle formalism to describe reference frames and relational observables. This approach, widely used in foundational studies of gauge theories (see e.g. Healey (2007); Weatherall (2016)), is primarily attributed to the work of Gomes (see, e.g., Gomes (2023a, b) for a rigorous treatment). Let $M$ represent the space of models $m$ of the theory. The space $M$ can be described as a principal bundle with $S$ as its structure group and $[M]:=\{[m]\mid m\in M\}$ as its base space, where $[m]$ identifies the equivalence class of models under the transformations of $\mathcal{S}$. In this formalism, selecting a reference frame involves defining a unique section map $\sigma:[m]\to\sigma([m])\in M$, which smoothly maps equivalence classes of models to individual models in the space . This corresponds to choosing a submanifold in the fibre bundle that intersects each fibre $\mathbb{F}_{m}:=\mathrm{pr}^{-1}([m])$ exactly once, with $\mathrm{pr}:m\to[m]$ being the projection map.
Choosing a reference frame can also be interpreted as selecting a specific gauge. In fact, the choice of a gauge is equivalent to the choice of a reference frame (see Dittrich (2007) for insights on the relationship between gauge-fixed observables and relational observables. See also Bamonti and Thébault (2024) for an example of such relationship in FLRW cosmology). In gauge theories with a principal fibre bundle structure, gauge-fixing determines a section through the fibre bundle. Given a symmetry group $S$, each fibre corresponds to a gauge orbit — the set of all configurations of a field that are related by gauge transformations, generated by the constraints of the theory. Specifically, there exists a one-to-one correspondence between each gauge orbit and the equivalence class of models under the symmetry $s\in S$. More concretely, for $m\in M$ and $s\in S$, the orbit $\mathcal{O}_{m}$ is defined as $\mathcal{O}_{m}=\{m_{s}\mid s\in S\}$, where $m_{s}$ denotes the model to which the transformation $s$ is applied. Due to the symmetry of the theory, the model space includes significant redundancy. The physical content of the theory is captured by a single representative from each gauge orbit, while other models within the same orbit are redundant, isomorphic copies. [^12] This does not imply that gauge freedom constitutes mere ”descriptive fluff” (see Earman (2004)). On the contrary, the correspondence between gauge-fixing and reference frame selection highlights the relational nature of physics. The additional degrees of freedom are meaningful as they represent the possible ways a system can form observables relative to another system. Isomorphic models provide ”handles” through which systems can couple; see Rovelli (2014) and Adlam (2024a). This redundancy is typically resolved by fixing a gauge. Similarly, in GR, one fixes a reference frame through a coordinate gauge condition (see Bamonti (2023); Gomes (2024b)).
@@ 2 The Fibre Bundle Formalism: A Gauge Perspective on Reference Frames
where $f_{\Re_{g}}$ acts as the equivalent of a section map within the fibre-bundle framework. Specifically, this diffeomorphism $f_{\Re_{g}}:g_{ab}\to f_{\Re_{g}}^{*}g_{ab}$ serves as a projection operator, uniquely mapping any element of a given fibre to the section’s image. It is also called the projection operator for the section. [^18] The projection corresponding to the choice of a reference frame can be written down in terms of coordinate charts. For instance, in GR two common gauge-fixings are the De Donder gauge in the Lagrangian sector which corresponds to the condition $F(g)=\partial_{\mu}(g^{\mu\nu}\sqrt{g})=0$, and the CMC gauge in the Hamiltonian sector: $F(h^{ij},\pi_{ij})=h^{ij}\pi_{ij}=const$. The former corresponds to the use of coordinates that satisfy a relativistic wave equation $\Box x^{\mu}=0$. The latter, selects global simultaneity homogeneous 3-hypersurfaces $\Sigma_{\tau}$, parametrised by a universal time $\tau$. Notice that both are only partial gauge-fixings. More precisely, $f_{\Re_{g}}$ is the embedding map, acting within a fibre, from the fibre bundle manifold of models $Lor(\mathcal{M})$ to the image of the section map, and it is characterised by the auxiliary condition $F_{\Re_{g}}(g_{ab})=0$. The constraint $F_{\Re_{g}}(g_{ab})=0$ defines a ‘level surface’ of the section map along the fibres, effectively making the choice of a reference frame (or section) analogous to a gauge-fixing procedure. See Figure 1.191919 According to the condition (3), the choice of a reference frame, seen as the choice of a section, is based on the imposition of a set of conditions that a models must satisfy. This procedure is analogous to what Gauss (1902) proposed in order to describe a surface embedded in an ambient space not from an external point of view, i.e. using the coordinates of the ambient space, but ‘standing on the surface itself’. Such an embedded surface is intrinsically describable using some parametric equations. For example, a generic $n$-dimensional surface $\Sigma\subset N$, embedded in a generic $(n+1)$-dimensional Euclidean ambient space $N$ (characterised by $n+1$ coordinates $x^{i}$), can be described by $n$ parameters $u^{\alpha}$ with parametric equations of the type: $x^{i}(u^{\alpha})=0$. In our case, such equations take the form $F_{\mathcal{R}_{g}}(g_{ab})=0$, with $F_{\mathcal{R}_{g}}\in C^{\infty}(\mathcal{M})$ being a smooth and regular function.  [^20] Given a general symmetry group, its gauge orbits are in general not one-dimensional. The representation of figure 1 is faithful only for one-dimensional groups, whose action can be depicted in a one-to-one manner along the one-dimensional orbits.